OpenStudy (anonymous):

What is the result of isolating y2 in the equation below? (x - 2)2 + y2 = 64 A. y2 = -x2 + 4x + 4 B. y2 = 64 - x2 C. y2 = x2 + 4x + 60 D. y2 = -x2 + 4x + 60

OpenStudy (photon336):

You can re-write it like this \[y^{2}+(x-2)^{2} = 64 \]

OpenStudy (anonymous):

B?

OpenStudy (photon336):

We need to expand (x-2)^2 first too. \[(x-2)^{2} = x^{2}-4x+4 \] \[y^{2}+x^{2}-4x+4 = 64\]

OpenStudy (anonymous):

where are the 4´s coming from

OpenStudy (photon336):

we needed to expand (x-2)^2 that's why

OpenStudy (anonymous):

ok how do u expand it

OpenStudy (photon336):

\[(x-2)(x-2) \]

OpenStudy (anonymous):

what is x

OpenStudy (photon336):

you foil it

OpenStudy (anonymous):

when i put that in i got -79200

OpenStudy (photon336):

\[x^{2}-4x+4 = (x-2)(x-2)\]

OpenStudy (anonymous):

i still dont know where the 4x+4 came from

OpenStudy (photon336):

this is what we end up with \[y^{2}+x^{2}-4x+4 = 64\]

OpenStudy (photon336):